OpenStudy (anonymous):

Derivative (Quotient Rule) f(x)= x^2/((squared root x)-3)

OpenStudy (anonymous):

I got it to here: \[\frac{ \frac{ 3 }{ 2 }x ^{3/2}-6x }{ (x ^{1/2}-3)^2 }\] I'm not sure if it's right

OpenStudy (anonymous):

[2x* (sqrt(x)-3)-x^2*(1/2(x)^(-1/2)]/(sqrt(x)-3)^2

OpenStudy (callisto):

\[f(x) = \frac{x^2}{x^\frac{1}{2}-3}\]\[\text{Let } g(x) = x^2, h(x) = x^\frac{1}{2}-3\] What are g'(x) and h'(x)?

OpenStudy (anonymous):

g'(x) = 2x h'(x)=1/2x^-1/2

OpenStudy (callisto):

Yup! \[f'(x) = \frac{h(x)g'(x) - g(x)h'(x)}{[h(x)]^2} = \frac{(x^\frac{1}{2}-3)(2x)-x^2(\frac{1}{2}x^{-\frac{1}{2}})}{(x^\frac{1}{2}-3)^2}\]

OpenStudy (callisto):

\[= \frac{2x^\frac{3}{2}-6x-\frac{1}{2}x^{\frac{3}{2}}}{(x^\frac{1}{2}-3)^2} = \frac{\frac{3}{2}x^\frac{3}{2}-6x}{(x^\frac{1}{2}-3)^2} \] Your answer looks good to me.

OpenStudy (anonymous):

but at the back of the textbook it says: \[f'(x)=\frac{ 3x \sqrt{x}-12x }{ 2(\sqrt{x}-3)^2 }\]

OpenStudy (callisto):

LOL, they are just the same! \[\frac{\frac{3}{2}x^\frac{3}{2}-6x}{(x^\frac{1}{2}-3)^2} = \frac{\frac{1}{2}(3x^\frac{3}{2}-12x)}{(x^\frac{1}{2}-3)^2} = \frac{3x^\frac{3}{2}-12x}{2(x^\frac{1}{2}-3)^2} =\frac{3x(x^\frac{1}{2})-12x}{2(x^\frac{1}{2}-3)^2} = \frac{3x\sqrt x-12x}{2(x^\frac{1}{2}-3)^2}\]

OpenStudy (anonymous):

Wow! Did not know that. Thanks for writing that for me! :)

OpenStudy (callisto):

Welcome :)